# A mesoscopic model for the rheology of dilute and semidilute solutions of wormlike micelles

### v1

@@ III.3.3 Length distributions and the fraction of ring wormlike micelles
Table 3: The quantity $\gamma_{\text{eff}}^{*}$, determined from $\gamma_{\text{eff}}^{*}=\mu\bar{L}$, matches the values obtained from fitting the scaled length distribution of linear chains. In the dilute regime ($c/c_{\text{pw}}^{*}=0.2$), $\gamma_{\text{eff}}^{*}\approx\gamma_{\text{eff}}=1.162$, and in the semidilute regime ($c/c_{\text{pw}}^{*}=1$), $\gamma_{\text{eff}}^{*}\approx\gamma_{\text{eff}}=1$.
$\epsilon_{st}$ $c/c^{*}_{\text{pw}}=0.2$ $c/c^{*}_{\text{pw}}=1$
@@
An alternative way to examine the influence of rings is to calculate the number of free stickers in solution, $\Upsilon$, in the presence and absence of rings (see Fig. 15), which can be calculated from,
where, $N_{\text{st}}^{\text{bound}}=N_{\text{st,lin}}^{\text{bound}}+N_{\text{st,R}}^{\text{bound}}\,,\>\text{with}\>\ N_{\text{st,lin}}^{\text{bound}}=\sum_{L_{\text{min}}}^{L_{\text{max}}}\,\left(2\,m_{\text{pw,lin}}^{\text{L}}-2\right)n^{\text{lin}}_{\text{L}}(L)$. The subtraction of $2$ in the definition of $N_{\text{st,lin}}^{\text{bound}}$ accounts for the fact that the end beads in linear chains remain unbound. Clearly, when the possibility of ring formation is permitted in the model, there are far fewer free stickers in solution, as the stickers at the end of linear chains combine and lead to the formation of rings. At high sticker energies, when the ends of even single persistent worms combine to form rings, stickers remain bound as the concentration increases. In both cases of rings and no rings, as the fraction of linear wormlike micelles grows with increasing concentration, the fraction of free stickers that are at ends of the chains decreases monotonically.
III.3.4 The radius of gyration for linear wormlike micelles
@@
It may be recalled that the overlap concentration $c^{*}$, at a fixed sticker energy $\epsilon_{st}$, was defined as the effective concentration at which the overlap length $L^{*}$ is equal to the mean length $\bar{L}$ (see Fig. 4). One can calculate the number of effective monomers $N^{*}$ that correspond to the overlap length $L^{*}$ from $N^{*}=(L^{*}/b)+1$. When estimated in this manner for $\epsilon_{st}=6$, it turns out that $N^{*}=N_{\mathrm{L,lin}}^{\mathrm{eff^{*}}}$, as determined in Fig. 16 (c). Thus, at the overlap concentration, the correlation blob is just equal in size to the size of the swollen coil conformation adopted by a linear wormlike micelle of overlap length $L^{*}$, reinforcing the validity of determining the overlap concentration by the procedure described in Sec. II.5. By fitting data at various values of $N^{*}$ determined by varying $\epsilon_{st}$ at $c^{\mathrm{eff}}/c^{*}=1$, it is found that $R_{g_{L}}^{*}(\epsilon_{st})=B\left(N^{*}\right)^{0.59}$, where $B\approx 1.41$ is a fitting parameter.
As defined earlier in Eq. (27), the alternative definition of the radius of gyration used here is $R_{g}$, which is the mean size of all linear wormlike micelles in the solution, regardless of their length. The square of this radius of gyration is displayed in Figure 17 (a) in the dilute regime, as a function of $\bar{N}$, which is the number of beads in a linear wormlike micelle with length equal to the mean length $\bar{L}$. Each of the data points in the plot correspond to simulation results obtained at values of $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ in Table 2 that lead to values of $c^{\text{eff}}/c^{*}<1$, with $\bar{L}$ evaluated at these particular values of concentration and sticker energy. The number of beads is calculated from $\bar{N}=(\bar{L}/b)+1$. Remarkably, the dependence of $R_{g}^{2}$ on $\bar{N}$ appears to be identical to that of ${R^{*}_{g_{L}}}^{\!\!\!2}$ on $N^{*}$, i.e., $R_{g}=B\left(\bar{N}\right)^{0.59}$, implying one can write $R_{g}/R^{*}_{g_{L}}=\left(\bar{N}/N^{*}\right)^{0.59}$, regardless of the value of $\epsilon_{st}$. Indeed, when the ratio $(R_{g}^{2}/{R^{*}_{g_{L}}}^{\!\!\!2})$ is plotted as a function of the ratio $(\bar{N}/N^{*})$ in both the dilute and semidilute regimes, as displayed in Fig. 17 (b), data for various values of sticker energy $\epsilon_{st}$ collapse on to a master plot, that exhibits a sharp transition at $\bar{N}/N^{*}=1$ from one power law regime with slope $1.18\pm 0.02$ to another power law regime with slope $1.0\pm 0.03$. This suggests that one could write $R_{g}=R^{*}_{g_{L}}\left(\bar{N}/N^{*}\right)^{\nu}$, where $\nu$ is the Flory exponent with value 0.59 in the dilute regime and 0.5 in the semidilute regime. Such a representation is consistent with the picture that at any sticker energy $\epsilon_{st}$ in the semidilute regime, the mean size of all the wormlike micelles in a simulation box is equivalent to that of a linear wormlike micelle of mean length $\bar{L}$, whose conformation is a random walk of correlation blobs, where $R^{*}_{g_{L}}$ is the size of the correlation blob at the overlap concentration, and $(\bar{N}/N^{*})$ is the number of such correlation blobs in the chain, while in the dilute regime where $\bar{N}$ is less than $N^{*}$, the mean size of all the wormlike micelles is equivalent to that of a linear wormlike micelle of mean length $\bar{L}$, whose conformation is a self-avoiding walk.
@@ III.4.3 Role of hydrodynamic interactions
Figure 22: Illustration of Rouse ($h^{*}=0$) and Zimm ($h^{*}=0.2$) dynamics in wormlike micelle solutions with changing concentration in the dilute and semidilute regimes, at a fixed sticker energy $\epsilon_{st}=7$. The first column presents the storage modulus $G^{\prime}$ at three concentrations: (a) $c^{\mathrm{eff}}/c^{*}=0.51$, (b) $c^{\mathrm{eff}}/c^{*}=4$, and (c) $c^{\mathrm{eff}}/c^{*}=6.50$, while the second column (d) to (f), displays the loss modulus $G^{\prime\prime}$ at the same three concentrations.
Fig. 22 displays plots of $G^{\prime}$ and $G^{\prime\prime}$ as a function of frequency at three different concentrations in the dilute and semidilute regimes, for both the cases where hydrodynamic interactions are switched on or switched off. The first column [(a) to (c)] presents the storage modulus $G^{\prime}$, while the second column [(d) to (f)] displays the loss modulus $G^{\prime\prime}$. The concentration is increased systematically, with $c^{\mathrm{eff}}/c^{*}=0.51$ in the first row, $c^{\mathrm{eff}}/c^{*}=4$ in the second row, and $c^{\mathrm{eff}}/c^{*}=6.50$ in the third row. As mentioned earlier, the range of frequencies in the intermediate frequency regime where a power law slope in $G^{\prime}$ is observed is limited because of the appearance of signatures of various molecular phenomena in this regime. This is even more so in the case of $G^{\prime\prime}$. Nevertheless, from the limited range of frequencies where data can be fitted to a power law, it can be seen from the inset to Fig. 22 (a) that the scaling with frequency is Rouse like ($\sim\omega^{1/2}$) in the absence of hydrodynamic interactions, and Zimm like ($\sim\omega^{1/3\nu}$, with $\nu\approx 0.6$) when they are taken into account. With increasing concentration, Figs. 22 (a) to (c) indicate that the screening of hydrodynamic interactions sets in, and the difference in the slopes of the power law regime between the two cases decreases, until it vanishes by a concentration of $c^{\mathrm{eff}}/c^{*}=6.50$, at which point only Rouse scaling is observed. Interestingly, a difference in the curves with and without hydrodynamic interactions persists for both $G^{\prime}$ and $G^{\prime\prime}$ at the high end of frequencies in regime II (before the crossover into regime III), regardless of the increase in concentration.
@@ IV Conclusions
    Simulation data for the ratio of the mean length of wormlike micelles to the overlap length, $\bar{L}/L^{*}$, as a function of the scaled concentration $c^{\mathrm{eff}}/c^{*}$, for different sticker energies $\epsilon_{st}$, collapses onto a universal master curve obeying $\bar{L}/L^{*}=\left(c^{\text{eff }}/c^{*}\right)^{\alpha}$, with $\alpha=0.46$ in the dilute limit and 0.6 in semidilute limit, with a sharp crossover from the dilute to the semidilute regime occuring at $\bar{L}/L^{*}=c^{\mathrm{eff}}/c^{*}=1$ (see Fig. 9). The influence of sticker energy is captured through the overlap length $L^{*}$ and the overlap concentration $c^{*}$, both of which vary exponentially with $\epsilon_{st}$ (see Fig. 10). The presence or absence of rings does not impact the scaling of linear chains, provided that the effective monomer concentration, $c^{\mathrm{eff}}$, is computed only for linear wormlike micelles. Scaling behaviour is not observed at low sticker energies when $c^{\mathrm{eff}}/c^{*}<0.2$, or if the linear micelles are too short, with $\bar{L}<20)$.
@@
    The dynamic moduli of wormlike micellar solutions exhibit many additional features in the intermediate frequency regime compared to those observed for monodisperse homopolymer solutions. The molecular origin of these features, and their connection to the timescales of microscopic phenomena, will be explored in future studies. The longest relaxation time of polydisperse wormlike micelles, $\tau_{G^{\prime}=G^{\prime\prime}}$, is found to be identical to that of a monodisperse homopolymer solution, provided that the homopolymer chain length and scaled solution concentration are equal to the mean length of the linear wormlike micelles, and the scaled effective concentration of the wormlike micellar solution, respectively (see Figs. 20 and 21).

### v2

@@ III.3.3 Length distributions and the fraction of ring wormlike micelles
Table 2: Values of parameters in Eq. (42) determined by fitting simulation data and the quantity $\gamma_{\text{eff}}^{*}$, determined from $\gamma_{\text{eff}}^{*}=\mu\bar{L}$. The values of $\gamma_{\text{eff}}^{*}$ match the values obtained from fitting the scaled length distribution of linear chains. In the dilute regime ($c/c_{\text{pw}}^{*}=0.2$), $\gamma_{\text{eff}}^{*}\approx\gamma_{\text{eff}}=1.162$, and in the semidilute regime ($c/c_{\text{pw}}^{*}=1$), $\gamma_{\text{eff}}^{*}\approx\gamma_{\text{eff}}=1$.
$\epsilon_{st}$ $c^{*}_{\text{pw}}=0.2 c/c$$c^{*}_{\text{pw}}=1$
@@
An alternative way to examine the influence of rings is to calculate the number of free stickers in solution, $\Upsilon$, in the presence and absence of rings, which can be calculated from,
where, $N_{\text{st}}^{\text{bound}}=N_{\text{st,lin}}^{\text{bound}}+N_{\text{st,R}}^{\text{bound}}\,,\>\text{with}\>\ N_{\text{st,lin}}^{\text{bound}}=\sum_{L_{\text{min}}}^{L_{\text{max}}}\,\left(2\,m_{\text{pw,lin}}^{\text{L}}-2\right)n^{\text{lin}}_{\text{L}}(L)$. The subtraction of $2$ in the definition of $N_{\text{st,lin}}^{\text{bound}}$ accounts for the fact that the end beads in linear chains remain unbound. Clearly, when the possibility of ring formation is permitted in the model, there are far fewer free stickers in solution, as the stickers at the end of linear chains combine and lead to the formation of rings. At high sticker energies, when the ends of even single persistent worms combine to form rings, stickers remain bound as the concentration increases. In both cases of rings and no rings, as the fraction of linear wormlike micelles grows with increasing concentration, the fraction of free stickers that are at ends of the chains decreases monotonically (Fig. 17).
III.3.4 The radius of gyration for linear wormlike micelles
@@
It may be recalled that the overlap concentration $c^{*}$, at a fixed sticker energy $\epsilon_{st}$, was defined as the effective concentration at which the overlap length $L^{*}$ is equal to the mean length $\bar{L}$ (see Fig. 5). One can calculate the number of effective monomers $N^{*}$ that correspond to the overlap length $L^{*}$ from $N^{*}=(L^{*}/b)+1$. When estimated in this manner for $\epsilon_{st}=6$, it turns out that $N^{*}=N_{\mathrm{L,lin}}^{\mathrm{eff^{*}}}$, as determined in Fig. 18 (c) above. Thus, at the overlap concentration, the correlation blob is just equal in size to the size of the swollen coil conformation adopted by a linear wormlike micelle of overlap length $L^{*}$, reinforcing the validity of determining the overlap concentration by the procedure described in Sec. II.5. By fitting data at various values of $N^{*}$ determined by varying $\epsilon_{st}$ at $c^{\mathrm{eff}}/c^{*}=1$, it is found that $R_{g_{L}}^{*}(\epsilon_{st})=B\left(N^{*}\right)^{0.59}$, where $B\approx 1.41$ is a fitting parameter.
@@
As defined earlier in Eq. (27), the alternative definition of the radius of gyration used here is $R_{g}$, which is the mean size of all linear wormlike micelles in the solution, regardless of their length. The square of this radius of gyration is displayed in Figure 19 (a) in the dilute regime, as a function of $\bar{N}$, which is the number of beads in a linear wormlike micelle with length equal to the mean length $\bar{L}$. Each of the data points in the plot correspond to simulation results obtained at values of $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ in Fig. 8 that lead to values of $c^{\text{eff}}/c^{*}<1$, with $\bar{L}$ evaluated at these particular values of concentration and sticker energy. The number of beads is calculated from $\bar{N}=(\bar{L}/b)+1$. Remarkably, the dependence of $R_{g}^{2}$ on $\bar{N}$ appears to be identical to that of ${R^{*}_{g_{L}}}^{\!\!\!2}$ on $N^{*}$, i.e., $R_{g}=B\left(\bar{N}\right)^{0.59}$, implying one can write $R_{g}/R^{*}_{g_{L}}=\left(\bar{N}/N^{*}\right)^{0.59}$, regardless of the value of $\epsilon_{st}$. Indeed, when the ratio $(R_{g}^{2}/{R^{*}_{g_{L}}}^{\!\!\!2})$ is plotted as a function of the ratio $(\bar{N}/N^{*})$ in both the dilute and semidilute regimes, as displayed in Fig. 19 (b), data for various values of sticker energy $\epsilon_{st}$ collapse on to a master plot, that exhibits a sharp transition at $\bar{N}/N^{*}=1$ from one power law regime with slope $1.18\pm 0.02$ to another power law regime with slope $1.0\pm 0.03$. This suggests that one could write $R_{g}=R^{*}_{g_{L}}\left(\bar{N}/N^{*}\right)^{\nu}$, where $\nu$ is the Flory exponent with value 0.59 in the dilute regime and 0.5 in the semidilute regime. Such a representation is consistent with the picture that at any sticker energy $\epsilon_{st}$ in the semidilute regime, the mean size of all the wormlike micelles in a simulation box is equivalent to that of a linear wormlike micelle of mean length $\bar{L}$, whose conformation is a random walk of correlation blobs, where $R^{*}_{g_{L}}$ is the size of the correlation blob at the overlap concentration, and $(\bar{N}/N^{*})$ is the number of such correlation blobs in the chain, while in the dilute regime where $\bar{N}$ is less than $N^{*}$, the mean size of all the wormlike micelles is equivalent to that of a linear wormlike micelle of mean length $\bar{L}$, whose conformation is a self-avoiding walk.
@@ III.4.3 Role of hydrodynamic interactions
Figure 24: Illustration of Rouse ($h^{*}=0$) and Zimm ($h^{*}=0.2$) dynamics in wormlike micelle solutions with changing concentration in the dilute and semidilute regimes, at a fixed sticker energy $\epsilon_{st}=7$. The first column presents the storage modulus $G^{\prime}$ at three concentrations: (a) $c^{\mathrm{eff}}/c^{*}=0.51$, (b) $c^{\mathrm{eff}}/c^{*}=4$, and (c) $c^{\mathrm{eff}}/c^{*}=6.50$, while the second column (d) to (f), displays the loss modulus $G^{\prime\prime}$ at the same three concentrations. In all these cases, the error bars in the simulation data are smaller than the symbol size.
Figure 24 displays plots of $G^{\prime}$ and $G^{\prime\prime}$ as a function of frequency at three different concentrations in the dilute and semidilute regimes, for both the cases where hydrodynamic interactions are switched on or switched off. The first column [(a) to (c)] presents the storage modulus $G^{\prime}$, while the second column [(d) to (f)] displays the loss modulus $G^{\prime\prime}$. The concentration is increased systematically, with $c^{\mathrm{eff}}/c^{*}=0.51$ in the first row, $c^{\mathrm{eff}}/c^{*}=4$ in the second row, and $c^{\mathrm{eff}}/c^{*}=6.50$ in the third row. As mentioned earlier, the range of frequencies in the intermediate frequency regime where a power law slope in $G^{\prime}$ is observed is limited because of the appearance of signatures of various molecular phenomena in this regime. This is even more so in the case of $G^{\prime\prime}$. Nevertheless, from the limited range of frequencies where data can be fitted to a power law, it can be seen from the inset to Fig. 24 (a) that the scaling with frequency is Rouse like ($\sim\omega^{1/2}$) in the absence of hydrodynamic interactions, and Zimm like ($\sim\omega^{1/3\nu}$, with $\nu\approx 0.6$) when they are taken into account. With increasing concentration, Figs. 24 (a) to (c) indicate that the screening of hydrodynamic interactions sets in, and the difference in the slopes of the power law regime between the two cases decreases, until it vanishes by a concentration of $c^{\mathrm{eff}}/c^{*}=6.50$, at which point only Rouse scaling is observed. Interestingly, a difference in the curves with and without hydrodynamic interactions persists for both $G^{\prime}$ and $G^{\prime\prime}$ at the high end of frequencies in regime II (before the crossover into regime III), regardless of the increase in concentration.
@@ IV Conclusions
    Simulation data for the ratio of the mean length of wormlike micelles to the overlap length, $\bar{L}/L^{*}$, as a function of the scaled concentration $c^{\mathrm{eff}}/c^{*}$, for different sticker energies $\epsilon_{st}$, collapse onto a universal master curve obeying $\bar{L}/L^{*}=\left(c^{\text{eff }}/c^{*}\right)^{\alpha}$, with $\alpha=0.46$ in the dilute limit and 0.6 in semidilute limit. A crossover occurs in the concentration range $0.7\lesssim c^{\mathrm{eff}}/c^{*}\lesssim 1$ (see Fig. 11). The influence of sticker energy is captured through the overlap length $L^{*}$ and the overlap concentration $c^{*}$, both of which vary exponentially with $\epsilon_{st}$ (see Fig. 12). The presence or absence of rings does not impact the scaling of linear chains, provided that the effective monomer concentration, $c^{\mathrm{eff}}$, is computed only for linear wormlike micelles. Scaling behaviour is not observed at low sticker energies when $c^{\mathrm{eff}}/c^{*}<0.2$, or if the linear micelles are too short, with $\bar{L}<20)$.
@@
    The dynamic moduli of wormlike micellar solutions exhibit many additional features in the intermediate frequency regime compared to those observed for monodisperse homopolymer solutions. The molecular origin of these features, and their connection to the timescales of microscopic phenomena, will be explored in future studies. The longest relaxation time of polydisperse wormlike micelles, $\tau_{G^{\prime}=G^{\prime\prime}}$, is found to be identical to that of a monodisperse homopolymer solution, provided that the homopolymer chain length and scaled solution concentration are equal to the mean length of the linear wormlike micelles, and the scaled effective concentration of the wormlike micellar solution, respectively (see Figs. 22 and 23).